OpenStudy (anonymous):

65% of 80 is what number? 25 52 81.3 123

OpenStudy (anonymous):

ok divide 80 into 100

OpenStudy (anonymous):

\[80/100=0.8\]

OpenStudy (anonymous):

\[0.8*65=52\]

OpenStudy (anonymous):

Okay thank you!!!!

OpenStudy (anonymous):

Do that or the easier and quicker way, just do 80 * 0.65 which equals 52.

OpenStudy (anonymous):

so the answer is \[B\]

OpenStudy (anonymous):

\[\frac{ x }{ 80 } = \frac{ 65 }{ 100 }\] \[65*80 = 5200/100x\] x=52

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

48 is 35% of what number?

OpenStudy (anonymous):

48 goes into 100 how many times? @tatumrocks

OpenStudy (anonymous):

48 is 35 percent of 137.14286